### v3

# InterIntent: Investigating Social Intelligence of LLMs via Intention Understanding in an Interactive Game Context

Abstract
Large language models (LLMs) have demonstrated the potential to mimic human social intelligence. However, most studies focus on simplistic and static self-report or performance-based tests, which limits the depth and validity of the analysis. In this paper, we developed a novel framework, InterIntent, to assess LLMs’ social intelligence by mapping their ability to understand and manage intentions in a game setting. We focus on four dimensions of social intelligence: situational awareness, self-regulation, self-awareness, and theory of mind. Each dimension is linked to a specific game task: intention selection, intention following, intention summarization, and intention guessing. Our findings indicate that while LLMs exhibit high proficiency in selecting intentions, achieving an accuracy of 88%, their ability to infer the intentions of others is significantly weaker, trailing human performance by 20%. Additionally, game performance correlates with intention understanding, highlighting the importance of the four components towards success in this game. These findings underline the crucial role of intention understanding in evaluating LLMs’ social intelligence and highlight the potential of using social deduction games as a complex testbed to enhance LLM evaluation. InterIntent contributes a structured approach to bridging the evaluation gap in social intelligence within multiplayer games. [^1] Code is available at https://github.com/uscnlp-lime/Inter-Intent
InterIntent: Investigating Social Intelligence of LLMs via
Intention Understanding in an Interactive Game Context
Ziyi Liuthanks:   Equal contribution.    Abhishek Anand     Pei Zhou    Jen-tse Huang   Jieyu Zhao University of Southern California {zliu2803, anandabh, peiz, jh_116, jieyuz}@usc.edu
1 Introduction
The growing intelligence of large language models (LLMs) has facilitated diverse research on their capability to mimic human social intelligence (Ziems et al., 2024; Dubova, 2022; Gweon et al., 2023; Huang et al., 2023; Sap et al., 2022; Wu et al., 2023). Social intelligence, the ability to understand and manage one’s own and others’ actions and to act wisely in social relations Thorndike (1920), usually includes four key components Silvera et al. (2001): (1) Situational Awareness: The perception and comprehension of the elements in the environment Endsley (1995). (2) Self-Regulation: The process of guiding one’s own thoughts, behaviors, and feelings to reach goals Bandura (1991). (3) Self-Awareness: The understanding of one’s own character, feelings, motives, and desires Gallup et al. (2003). (4) Theory of Mind (ToM): The knowledge about others’ beliefs, intentions, and thoughts Baron-Cohen (1991).
Figure 1: Four dimensions to assess social intelligence in Avalon. We provide a dynamic and complex gaming context for evaluations. For situational awareness, we provide both positive and negative examples. In the negative example, since the previous quest was successful and no player was in a failed quest, the intention is inappropriate as there was no failed quest. For self-regulation, we require models to provide explicit information rather than repeating the intentions. Intentions are in bold within the contexts.
Despite adaptations for interactive social environments Zhou et al. (2023c, a); Park et al. (2023), most existing evaluations on LLMs Le et al. (2019); Shapira et al. (2023) focus on straightforward and static daily scenarios without explicitly defined goals for them. To enhance the contexts with complexity and dynamics with clear goals for LLMs, we turn our attention to social deduction games. These games require social interactions among players, providing a more diverse test bed for LLM evaluation Qiao et al. (2023). Existing studies have successfully utilized various games to analyze LLMs’ social behaviors, such as deception and confrontation Liang et al. (2023); Ibraheem et al. (2022); Mansbach et al. (2021); Meta et al. (2022); O’Gara (2023); Xu et al. (2023a, b); Wu et al. (2024a). However, these studies often focus on ad-hoc post-analysis of game performance and overlook systematic measurement of social intelligence, limiting the comprehensive understanding of LLMs’ capabilities in social environments.
In this paper, we leverage one of the representatives in social deduction games, Avalon Light et al. (2023a); Wang et al. (2023), as the context for LLM evaluation. Avalon is a social deduction game that relies on conversation, making it an ideal test bed with its goal-driven objectives and complex mechanisms, facilitating interactions among LLMs in a grounded environment. We systematically evaluate LLMs’ social intelligence across the four aforementioned dimensions. Specifically, we propose to focus on players’ Intentions Malle et al. (2001) at various phases of the game, since comprehension of one’s and others’ intentions plays a critical role in games Goodie et al. (2012). While previous research has explored intention detection (Zhang et al., 2020; Casanueva et al., 2020), the field remains largely underexplored in game contexts. As shown in Figure 1, we design four tests: (1) Intention Selection, (2) Intention Following, (3) Intention Summarization, (4) Intention Guessing, respectively for the four social intelligence dimensions.
Figure 2: The Avalon game process for one round. Left: the entire game pipeline. Right: the procedure for generating a single player’s speech.
We first develop a novel Avalon gaming framework enhanced with an intention-guided mechanism, drawing inspiration from previous studies Lan et al. (2023); Shi et al. (2023); Wang et al. (2023); Light et al. (2023b). Building on this, we introduced InterIntent, which is designed to dynamically generate contexts for evaluating social intelligence through LLMs’ intention understanding. With human annotation, our results reveal that GPT-3.5 and GPT-4 can select intentions correctly, with accuracies of 87.5% and 88.8%, respectively. GPT-3.5 achieves a 69.5% $F_{1}$ score in intention summarization, with GPT-4 reaching 83.8%, surpassing human performance. Furthermore, we find a strong correlation between high levels of intention comprehension and superior game win rate, highlighting the importance of the intention understanding components in enhancing LLMs’ gaming proficiency. However, the models fall short in intention following, with only 47.5% and 64.8% of responses rated as “Good” in human evaluations, and underperform in intention guessing, trailing human performance by 20%.
Our contributions are summarized as follows:
  • We bridge the gap between the evaluation of LLMs’ social intelligence and LLM-based multiplayer social games.
  • We introduce InterIntent, a framework designed to systematically analyze social intelligence through players’ intention understanding within the Avalon game environment.
  • Extensive experiments on various LLMs reveal their underperformance in intention guessing against human performance.
2 Intention-Guided Avalon Framework
We begin by outlining the entire game pipeline before focusing on the mechanics of intention-guided gameplay, which includes the categorization and application of intentions throughout the game.
2.1 LLM-Based Avalon Game
Avalon is a social deduction game where players are either loyal or evil, aiming to succeed in or sabotage quests through strategy, persuasion, deduction, and deception. The primary goal is for loyal players to complete quests, while evil players aim to fail them. [^2] The detailed rules of Avalon can be found in https://avalon.fun/pdfs/rules.pdf. Leveraging the publicly available prompts proposed by Wang et al. (2023), which includes first-order and second-order reasoning before speaking, we build a framework allowing five to ten players in gameplay (we set the players number as five in this paper which includes 2 evil players and 3 loyal players). As shown in Figure 2, each game round consists of five principal components: summarizing previous rounds, discussing strategies, reconsidering the team proposal by the leader, voting, and executing the quest. Notably, we introduce an additional component, the leader’s reconsideration phase as a critical, previously unexplored component in Avalon literature.
2.2 Intention Categorization
Score Criterion Example
1 The content does not mention the intentions at all. I am a loyal player and I support this proposal.
[…]
Table 1: Criteria for annotating intention following (speaking), accompanied by examples. The intention used across examples is “Express concerns about a player from a failed quest team and suggest to not include them in the current team.” The context is “The team proposal is Player 1 and Player 2, with Player 1 being on a failed quest.” We consider the score of “including wrong context knowledge” higher than “copy and paste” because we focus on whether the intention following has an impact on the game.
To study intentions within Avalon gameplay systematically, we adopt the definition by Kennington (2022), viewing intention as a choice coupled with commitment. This perspective expands intention beyond a mere mental state to include a discernible commitment to act with purpose. In line with Wang et al. (2023) and Xu et al. (2023a), we identify seven categories of intentions in LLMs: Interrogation, Defense, Confrontation, Concealment, Deception, Persuasion, and Teamwork. Initially, we derive a list of intentions from our direct experiences with the game of Avalon. We then utilize GPT-3.5 to extend these initial categorizations. Human annotators subsequently refine this expanded list, removing any intentions deemed unreasonable or redundant, thus producing a concise set of relevant intentions. This refinement is iterative, resulting in the final enumeration detailed in Appendix Sec. B.1.
2.3 Intention-Guided Game Playing
We introduce an intention-guided gameplay mechanism that integrates our defined intentions within the game environment, offering two main benefits. First, it enhances the performance of LLMs by focusing on explicit intention discussions. Second, it facilitates the evaluation of the four social intelligence components.
As depicted in Figure 2, the discussion phase involves several key steps to facilitate intentional interaction among players. Initially, players use a first-order prompt for deductive reasoning about each player’s role and intentions. Subsequently, players select two or three intentions from a predefined set according to the current game context. This selection process guides players in generating their thoughts and statements based on the chosen intentions. Players then engage in second-order reasoning to evaluate how their statements might be interpreted by others, allowing them to reconsider and adjust their initial intentions. Finally, players express their refined thoughts and statements during the game’s discussion segment, enhanced by this iterative reflection process. Full prompts are shown in Appendix Sec. E.
3 Intention-Centric Evaluations
This section introduces our evaluations, designed around the concept of intentions, and discusses their correlation with the four components of social intelligence outlined in Sec. 1. Intention is fundamental to social intelligence, essential for effective communication Tomasello (2023), strategic influence, adaptation to dynamic social interactions Zelazo et al. (2023), and achieving desired outcomes by helping us interpret and predict behaviors Yott and Poulin-Dubois (2016). Our evaluation order corresponds to the game-playing process.
3.1 Situational Awareness: Intention Selection
Situational awareness is the ability to perceive environmental elements over time and space, understand their significance, and predict their future status Li et al. (2024a); Endsley (1995). In our study, we assess the situational awareness of LLMs in social contexts by evaluating their intention selection. LLMs are requested to select intentions based on summarization, first-order reasoning, and the ongoing dialogue in the current round. Ideally, LLMs should exhibit a sharp awareness of dynamic interactions, choosing intentions that are contextually appropriate rather than contradictory. We deem intentions unreasonable if they are inconsistent with:
  • Established facts—for instance, when all players have voted in the previous proposal, a player intends to question why someone did not vote.
  • Role profiles—such as an evil player forgetting their identity and intending to mistakenly support the loyal side.
  • Other intentions—like a player intends to play Merlin and Percival simultaneously.
We use binary indicators to evaluate the reasonableness of intentions, assigning 1 to reasonable intentions and 0 to unreasonable ones. Previous hallucinations, such as manipulated information, might influence the current player’s decisions. However, we maintain strict criteria, as we expect the player to be capable of recognizing these hallucinations.
3.2 Self-Regulation: Intention Following
Self-regulation involves guiding one’s own thoughts, behaviors, and feelings to achieve goals, thereby requiring individuals to contribute to their own motivation Fitzsimons and Finkel (2010). In our study, we evaluate the self-regulation abilities of LLMs by assessing their adherence to selected intentions, which includes two main perspectives: thinking (planning) and speaking (implementing), as depicted in Figure 2. Due to the abstract nature of the thinking phase, our criteria are lenient: we consider the LLM’s thought process as correct if it reflects the intended goal. While we do not demand informativeness in the thinking phase, we require validity and penalize scores for hallucinations or omission of intentions. In contrast, for the speaking phase, we require the models to execute actions informatively and without hallucinations for a response to be considered good. We employ a Likert scale to annotate outcomes for both phases, with scores ranging from 5 (completely correct) to 1 (completely incorrect), with 3 representing a borderline. Table 1 provides detailed criteria and examples for the annotation.
3.3 Self-Awareness: Intention Summarization
Self-awareness refers to an individual’s understanding of their character, emotions, motives, and desires Gallup et al. (2003). Our evaluation investigates the capability of LLMs to accurately identify their own intentions through analysis of their internal thought processes and speeches in the current round. This evaluation serves as the converse to intention following. LLMs are expected not only to execute intentions precisely but also to articulate their underlying motivations.
3.4 Theory of Mind: Intention Guessing
ToM involves understanding others by attributing mental states to them Kosinski (2023). Previous evaluations of LLMs’ ToM focused on scenarios where models are provided with complete contexts for interpreting characters’ mental states Gandhi et al. (2024). Our evaluation raises the complexity and challenge for LLMs by providing only limited information, simulating real-world conditions. Specifically, LLMs are required to deduce players’ intentions from their speeches alone, representing a more rigorous test of their capability to comprehend and anticipate mental states.
4 Experimental Settings and Results
We run 40 games using gpt-3.5-turbo-1106 and 5 games using gpt-4-1106, with details in Appendix Sec. A. Sample size statistics across all experiments are presented in Table 2, and game performance metrics are discussed in Appendix Sec. C.
Models GPT-3.5 GPT-4
Intention Selection 2,440 350
Intention Following 7,507 1,278
Intention Summarization 2,316 261
Intention Guessing 2,235 201
Table 2: Statistics for sample size over all experiments. Intention following includes thinking and speaking.
4.1 Human Annotation
Intention Selection and Following
Figure 3: Self-regulation results. The results show the percentage of each score over all data samples. Scores 1-5 are evaluation criteria (Table 1). Score 5 means the best while score 1 means the worst.
Figure 1 illustrates that intention selection and following necessitate extensive human annotation. We obtain a total of 6,316 intentions in the 40 games utilizing GPT-3.5, presenting a significant workload for annotators who are required to meticulously examine every thinking and each speaking result. Therefore, we first annotate intentions from five games to identify those with significant, either positively or negatively, impacts on game outcomes. These “impactful intentions” are then prioritized for further study. The methodology for selecting these intentions and the correlation between intention understanding and game performance is detailed in Appendix Sec. B.2. This refinement process isolates 2,440 intentions from GPT-3.5 and 350 intentions from GPT-4 for annotation (Table 2).
We recruit five computer science master’s students for annotation, each assigned seven to eight game contexts. [^3] Hourly payment for all annotators in this work is \$16. Unimpactful intentions within each context are masked to maintain focus. To calculate inter-rater reliability, two files are commonly assigned across all annotators. The Fleiss’ Kappa Fleiss (1971) is calculated pairwise, with averages derived across all pairs. For the intention following, we group the scores into two categories: 1$\sim$3 and 4$\sim$5. Scores of 5 represent a perfect response, while 4 indicates a nearly perfect response. Scores from 1 to 3 signify that the response contains significant flaws. The scores are shown in Table 3. Additional grouping results are provided in the Appendix Sec F.
Intention Selection Intention Following (Thinking) Intention Following (Speaking)
Fleiss’ Kappa $0.69_{\pm 0.08}$ $0.54_{\pm 0.06}$ $0.49_{\pm 0.06}$
Table 3: Inter-rater agreement score among human annotators. We show mean $\pm$ standard deviation over pairs of annotators.
Intention Summarization and Guessing
Evaluations of intention summarization and guessing are separated from the gameplay pipeline because as interactions progress, the clarity of LLMs’ outputs often deteriorates rather than improves. At the game’s end, we extract the contexts including roles, discussions, voting outcomes, and quest results, in a structured format. Notably, intention summarization and guessing are not critical to the game’s core mechanics. By separating these elements, we can still ensure the comprehensiveness of structured game information, as detailed in Appendix Sec. D.
Summarizing and guessing intentions do not require human annotation, as we can directly compare the outcomes with players’ earlier choices. However, evaluating the performance of LLMs against humans remains essential. Unlike the straightforward tasks of intention selection and following, which generally require the exclusion of incorrect information—a task in which humans excel—summarizing and guessing intentions demands more social intelligence, a more complex human capability. We conduct a user study with three master’s students who answer a total of 300 questions, 200 for GPT-3.5, and 100 for GPT-4.
(a) Intention Selection (Game Wise)
(b) Intention Following $\geq$3 (Game Wise)
(c) Intention Following $=$3 (Game Wise)
Figure 4: Correlation between Intention Selection/Following and game performance. We present the percentages of games where evil players are equally, better, or worse than loyal players. For example, in games won by loyal players in (a), their performance matches or exceeds that of evil players. We mark the performance differences between evil and loyal players in red, showing a greater gap in successful games/quests compared to failed ones.
4.2 Results: Intention Selection and Following
LLMs show a good understanding of the situation.
For intention selection, we calculate the number of reasonable intentions over all intentions as accuracy. The accuracies for GPT-3.5 and GPT-4 are 87.5% and 88.8% respectively, indicating that both models effectively understand the situation. We do not assess whether the selection is optimal due to its subjective nature. Our focus is solely on whether the models can capture key information and make accurate decisions based on that information. Instances of unreasonable intentions are primarily attributed to the models forgetting information during lengthy conversations or being influenced by hallucinations from other players.
LLMs can understand intentions abstractly but fall short in planning out intentions.
For intention following, as shown in Figure 3, GPT-4 generally outperforms GPT-3.5, in both analytical thinking and articulate speaking. Specifically in speech, GPT-4’s responses tend to be more informative, comprehensive, and detailed. However, there are instances where GPT-4 deviates from its intended focus, although their discussion still makes sense. This misalignment with the intentions selected occurs approximately 15% of the time. It also appears that adhering to intentions in speaking proves more challenging than in thinking due to stricter evaluation criteria. This discrepancy suggests that while the model can often conceptualize an appropriate strategy, translating these strategies into clear, actionable responses—particularly in verbal form—remains a complex task.
Figure 5: ToM results over rounds. We provide 200 data points for human results on the GPT-3.5 games, and since usually, games stop at round 4, the results from round 5 are not included.
Intention understanding correlates with game performance.
To better understand how intention selection and following influence game dynamics, we examine their correlation with both game win rates and quest win rates separately. For the correlation between intention selection and game performance, we select intentions with a following score greater than two. For intention following, we consider only reasonable intentions, converting Likert scale responses to binary scores using varying thresholds. We define $r=\frac{sum(scores)}{n_{teammate}}$ where the score is the binary score result from intention selection/following and $n_{teammate}$ is the number of teammates, two for the evil side and three for the loyal side. If $r_{evil}>r_{loyal}$, it suggests that the evil players outperformed the loyal players in a given game or round, and vice versa. We use games from GPT-3.5 for assessment, whose results are shown in Figure 4 and Figure 6. We observe that loyal players outperform evil players across all quests or games, even in those they ultimately lose. This discrepancy likely arises from the loyal players’ conditions under more constrained information, leading to their occasional losses despite having generally superior performance. When comparing failed and successful games or quests, the performance gap between loyal and evil players is notably larger in successful ones than in failed ones, indicating that loyal players must significantly outperform evil players to secure a win. Specifically, in Figure 4 (c), a threshold of 3 is applied, labeling scores of “3” as “1” and all others as “0.” This threshold can reflect instances of hallucination or incorrect context information. Hallucinations can significantly influence players, yet incorrect adherence to intentions may occasionally yield positive outcomes. For example, if a loyal player mistakenly attacks an evil player for an incorrect reason, other players may overlook the mistake and continue to support the action. Our results suggest a correlation between selecting and following intentions and game performance, especially for loyal players at a disadvantage. Improved selections and following can enhance overall game performance.
4.3 Results: Intention Summarization and Guessing
LLMs can achieve human-level performance in self-awareness but perform much worse than humans in ToM.
For both summarization and guessing, we report F1 scores. As shown in Table 4, GPT-3.5 performs below human levels but still achieves respectable scores in self-awareness, while GPT-4 exhibits performance comparable to that of humans. One key reason for this difference is that “thinking” content more explicitly outlines intentions, making it easier for models to extract. From the ToM results shown in Table 5, we can see both models struggle significantly compared to self-awareness or compared to human performance. The experimental setup involves asking GPT-3.5 and GPT-4 to predict each other’s intentions, a task differing from self-awareness where a model summarizes its own intentions. In ToM assessments, models need to understand and predict intentions from responses other than their own, hence the lower performance of GPT-3.5, especially when interpreting the more nuanced and lengthy responses of GPT-4. This complexity results in GPT-3.5 becoming worse when predicting GPT-4’s responses, while GPT-4 manages similar accuracy in guessing both its own and GPT-3.5’s intentions. Additionally, human participants generally score better with GPT-4, suggesting that longer contexts in GPT-4’s outputs help humans and occasional errors in GPT-3.5’s outputs can sometimes confound human judgment. While GPT-4 aligns more closely with human-like processing, there is still considerable potential for improvement in understanding and mimicking human cognitive processes.
Incorporating more context does not necessarily improve the model’s ToM ability.
We also evaluated the impact of information accumulation on the ability to guess intentions, examining how the inclusion of information from previous rounds affects guessing results. Our analysis shown in Figure 5 reveals that for GPT-3.5 game outputs, more information generally leads to improved guessing outcomes. In contrast, the guessing scores for GPT-4 game outputs fluctuate significantly, with notable declines in GPT-3.5’s performance. This could be attributed to GPT-4 game outputs typically providing more extensive context, which challenges GPT-3.5’s capacity to effectively summarize and extract relevant information. Meanwhile, GPT-4 demonstrates a consistent ability to handle increased data volumes. Interestingly, human participants maintain similar scores across rounds. This might stem from humans’ ability to directly infer intentions from immediate thinking and speaking cues, without heavily relying on the context. To improve ToM capability, including more context does not necessarily improve the performance but might introduce more complexity to LLMs.
Games GPT-3.5 GPT-4
GPT-3.5/4 $69.49$ $83.76$
Human $75.21_{\pm 0.97}$ $83.34_{\pm 2.72}$
Table 4: Self-awareness results. The results are GPT-3.5 and GPT-4 summarize their own intentions respectively. We also show the mean and standard deviation of human study results since we include 3 human annotators. We show $F_{1}$ score in this table.
5 Related Work
5.1 Social Intelligence in LLMs
Artificial social intelligence has long been a significant topic (Bainbridge et al., 1994; Kovač et al., 2023; Grossmann et al., 2023; McDonald and Pearson, 2019; Street, 2024; Li et al., 2024b) as it enables machines to understand and respond to human emotions and social cues, enhancing human-computer interaction and facilitating social science study (Ziems et al., 2024; Griffin et al., 2023; Dubova, 2022; Gweon et al., 2023; Zhang et al., 2024). Gweon et al. (2023) proposed a dataset to evaluate social reasoning and Nematzadeh et al. (2018) crafted a dataset following Sally-Anne experiment to access ToM ability. Research has demonstrated the emergence and limitations of ToM abilities in LLMs (Kosinski, 2023; Moghaddam and Honey, 2023; Sap et al., 2022; Wu et al., 2023). Although LLMs can understand human intent and choose correct actions in some tasks, their overall performance remains inferior to that of humans.
5.2 Intentions and Theory of Mind
ToM, commonly referred to as “mindreading,” is traditionally understood as the ability to assign mental states to oneself and others. This capability is used to interpret and predict actions (van Duijn et al., 2023). Despite the emergence of ToM in LLM, some papers argued that LLM learns spurious correlations and ToM ability falls with stress test (Shapira et al., 2023; Ullman, 2023; Sclar et al., 2023). Intentions, as the driving force behind ToM (Kennington, 2022), is important as the primary purpose of comprehending beliefs and desires is to understand the intentions of others (Zelazo et al., 2023). Since social intelligence and ToM can be seen as a society of individually simple “agents” (Zhuge et al., 2023), we focus on intent understanding as an “agent” to evaluate social intelligence.
Games GPT-3.5 GPT-4
GPT-3.5 $31.73$ $24.85$
GPT-4 $45.66$ $46.87$
Human $61.34_{\pm 2.75}$ $65.23_{\pm 5.03}$
Table 5: ToM results of GPT-3.5 and GPT-4, along with the mean $\pm$ standard deviation of human results. Different from self-awareness, two LLMs also guess each other’s intentions. For example, when GPT-4 guesses GPT-3.5 intentions, it achieves a $F_{1}$ score of 45.66.
5.3 Social Game Context
With the enhancement of LLM performance, more research shifts to social games that necessitate conversation and cooperation (Ibraheem et al., 2022; Mansbach et al., 2021; Meta et al., 2022; O’Gara, 2023; Park et al., 2023; Zhou et al., 2024; Wu et al., 2024b). Zhou et al. 2023b used reinforcement learning to incorporate intent and ToM in DnD games. Recent works focus on the two most popular board games Avalon Lan et al. (2023); Shi et al. (2023); Wang et al. (2023); Light et al. (2023b) and Werewolf Xu et al. (2023b, a), exploring social behaviors and strategic playing ability of conversational agents. Unlike previous work which aimed to improve the game performance, we focus on the in-depth social intelligence evaluation utilizing the Avalon game.
6 Conclusion
We introduce a novel framework for evaluating social intelligence within the Avalon context, focusing on intentions. We examine four critical dimensions: situational awareness, self-regulation, self-awareness, and ToM. Our findings indicate that models demonstrate a fundamental understanding of situations and self-awareness, but struggle with self-regulation and ToM, particularly when compared to human performance. Through this systematic approach, our study not only contributes to a deeper understanding of LLMs’ social intelligence but also establishes a methodological foundation for future research.
Limitations
This study includes several limitations. First, we evaluate the four essential components of social intelligence but leave several aspects unexplored. For instance, the capability of a model to modify its intentions post-reflection, i.e., the Self-Correction, could also be assessed using our framework. Future research should consider expanding this approach to include more nuanced facets of social intelligence. The second limitation lies in our annotation criteria, which are restricted to assessing the reasonableness of intentions to avoid subjectivity. However, dimensions such as creativity and ingenuity are equally important. Future studies should aim to measure these attributes to ensure that models not only perform tasks effectively but also demonstrate innovative behaviors indicative of genuine intelligence. The third limitation stems from the cost. Half of our evaluations rely on costly human annotation, and the use of GPT-4 incurs significant resource demands due to the extensive context involved. To make the evaluation process more efficient and cost-effective, we need to develop methods that allow simpler tasks to be managed by smaller models, which introduces additional challenges. The final limitation is that we did not use an open-source model, as it lacks the capability to effectively engage in social games.
Ethics Statement
In order to achieve a smooth human-AI interaction, we need to assess and improve LLMs’ intelligence at a social level. Particularly, social intelligence plays an important role in human’s daily life and it would be a potential risk if models cannot handle complex social scenarios well. As a first step to building LLMs with human-like social intelligence, our framework provides a testbed to measure the difference between LLMs and humans in social intelligence.
Appendix A Models and Paramters
We only test GPT-3.5 and GPT-4 in our work as those two are the most common models in the social game study. We also tried open-source model like LLaMA-2 7B and LLaMA-3.1 8B, however, LLaMA-2 could not follow the format instructions properly, as also observed in Wang et al. (2023). LLaMA-3.1 performs better than LLaMA-2 but still fails to follow the format after a few turns of conversations. We set the temperature as 0.8 for both GPT-3.5 and GPT-4. For intention summarization, we also tried sampling five responses and got the aggregated result by majority vote. The result on multiple responses is only 1% higher than the result on single responses. Therefore, we only generate a single response for self-awareness and ToM tasks. The cost of GPT-3.5 is approximately \$80 and \$25 for GPT-4.
Appendix B Intentions list
B.1 Creation of Intention Set
The intentions were created as a combination of domain research and prompting GPT-3.5 to select intention categories from the options listed in Table 9 and generate intentions based on the selected categories. For intention category selection we use the prompt shown in Table 7 and for intention generation, we use the prompt shown in Table 8. The final set of intentions is listed in Table 9.
B.2 Selecting Impactful Intention
For selecting a set of impactful intentions, we annotate five complete games played by GPT-3.5. Further, we compute the probability of winning the round given that the intention was selected by a player. To select impactful intentions, we only consider the set of intentions where the computed probability is greater than 0.7 (strong positive association) or it is less than 0.3 (strong negative association) and the intentions selected at least two times for the considered side. We depict the final set of sixteen impactful intentions in red in Table 9.
Appendix C Game Performance
We study the game performance from various levels and perspectives:
Win Rate
It is the percentage of games won by the loyal side or the evil side from the total number of games.
Quest Win Rate
It is the percentage of quests (game rounds) won by a particular side from the total quests played in all the games. For the loyal side, a successful quest is a quest win, whereas, for the evil team, a failed quest is a quest win.
Quest Engagement Rate
It is the percentage of quests (game rounds) where the player gets selected in the quest team from the total quests played in all the games. This metric is calculated for each role separately and averaged across.
Team Selection Accuracy
It is the percentage of quests in which the leader correctly selected team members aligned with the objective of their respective side (loyal or evil), out of the total number of quests the leader belonged to that side. Correct team selection for loyal and evil sides is defined as follows:
  • Loyal: The proposed team by the loyal player contains all loyal players.
  • Evil: The proposed team by the evil player contains at least one evil player.
Failure Vote Rate
It is the percentage of failure votes from the total votes cast by evil players when they are included in the quest team.
Team Proposal Change Rate
It is the percentage of game rounds where the leader changes the proposed team after discussion, which reflects the effectiveness of the discussion.
Merlin Assassination Rate
At the end of the game, Assassin has a chance to identify Merlin. If they identify Merlin, the leader of the loyal side, correctly, then evil players win. This rate is the percentage of games where Assassin correctly identifies Merlin at the end of games over the games where Assassin has a chance to assassin.
We record the entire game and extract intermediate steps, such as voting results, team proposals, and quest outcomes, to calculate game performance. Results in Table 6 indicate that loyal players are at a disadvantage with a notably low game win rate, which corroborates findings from Wang et al. (2023). While the quest win and engagement rates show a balanced performance between loyal and evil players, the evil side demonstrates superior accuracy in team selection. This observation is consistent with their strategic role in the game. The high rate of team proposal changes across all settings suggests a dynamic framework where players consider the discussions of others, an essential aspect of social games. Additionally, the high rate of Merlin assassinations significantly boosts the win rate for evil players. Our objective is not to enhance game performance per se, but to demonstrate that with an intention-guided mechanism, game performance is similar to existing studies, providing a robust testbed for studies in social intelligence.
Metrics GPT3.5 GPT4
Loyal Evil Loyal Evil
Win rate 18.0 82.0 40.0 60.0
Quest win rate 41.4 58.6 54.5 45.5
Quest engagement rate 53.6 43.5 59.1 38.6
Team selection accuracy 32.1 80.0 71.4 100.0
Failure vote rate - 79.9 - 86.7
Team proposal change rate 76.2 74.1 50.00 62.5
Merlin assassination rate - 41.7 - 33.3
Table 6: Game performance results shown in percentage. For the failure vote rate and Merlin assassination rate, there is no result for Loyal players as they are not allowed to vote failure or assassinate Merlin.
Appendix D Structured Context Prompts
For intention summarization and intention guessing, we use structured context prompts. The detailed prompt structures are shown in Table 10 and Table 11.
Appendix E Game Pipeline Prompts
Since we mostly refer to the prompt by Wang et al. (2023), we only present prompts different from theirs to reduce the redundancy. The prompts we use are shown in Figure 7 to Figure 13. For the first order, formulation contemplation, and refinement contemplation, we add an intention mechanism based on original prompts. We also add several questions in the quest action prompt to help evil players better choose action.
Appendix F Human Annotation Details
We show all the instructions we use in this section, from Figure 14 to Figure 23. For intention summarization and guessing, we also illustrate the whole intention lists shown in Table 9.
We experimented with several grouping methods to calculate agreement scores. When using the grouping of 1 $\sim$ 4 and 5, the Fleiss’ Kappa scores for "thinking" and "speaking" were 0.448 and 0.420, respectively. These scores are relatively lower, likely due to differing interpretations of informativeness among human annotators. For the grouping of 1 $\sim$ 2 and 3 $\sim$ 5, the agreement was low, as scores of 1 $\sim$ 2 accounted for only a small portion of the entire dataset. Since Fleiss’ Kappa is highly sensitive to disagreements within the 1 $\sim$ 2 group, even minor discrepancies resulted in an overall lower score.
Intention Category Selection Prompt
Intentions are used to guide your game playing and they should align with your game Goal(main intent) and your role. Your selected intentions should support your main intent. Now select multiple intention categories that are ideal for this round depending on your role and strategy from the options listed below. Also give explanation of why you chose those categories in 2-3 sentences as part of Think value. Your explanation won’t be shown to others. Let’s think step by step before making your decisions. Remember to select at least 3 or more intentions from the given options only. intention category options are: [[[[intentioncategory options]]]]
Table 7: Intentions category selection prompt.
Intention Generation Prompt
First answer the following questions: 1. What is your main goal(main intent) for the game? 2. What have you observed from discussions in this round? 3. What should be the ideal team composition for your side to win? 4. After analyzing the responses to previous questions, what do you need to do to win from here? Consider answers to your previous questions and your roles and characteristics and generate intentions that help you achieve your main goal (main intention). Please do not repeat intentions and ensure they are different from each other. Each intention should be 10-15 words long. Let’s think step by step before making your decisions. Remember to generate at least 3 intentions, with a minimum of one intention for each intention category you have selected below: [[[[intentioncategories selected by player]]]]
Table 8: Intentions generation prompt.
(a) Intention Selection (Round Wise)
(b) Intention Following $\geq$3 (Round Wise)
(c) Intention Following $=$3 (Round Wise)
Figure 6: Correlation between Intention Selection/Following and game performance. We present the percentages of games where evil players are equally, better, or worse than loyal players. For example, in games won by loyal players in (a), their performance matches or exceeds that of evil players. We mark the performance differences between evil and loyal players in red, showing a greater gap in successful games/quests compared to failed ones.
Category Intentions
Interrogation Question a player on why he didn’t vote for the last team
Question a player about why they changed their statements
Question the leader why they selected a particular player for the team
Question the leader why they didn’t put themselves on the quest team
Defense Defend the proposed team composition if it includes loyal players (for Merlin)
Defend yourself against allegations that you could be evil
Defend your teammate against allegations that they could be evil (for evil players)
Confrontation Share concerns about an evil player (for Merlin)
Express concerns about a player from a failed quest team and suggest to remove them from current team
Cast suspicion on innocent players (for evil players)
Counter the team proposal citing that you are good and not part of the team
Counter the team proposal citing that you think a player on the team is evil
Express disagreement and vote disagree only if you are not in the proposed team
Concealment Stay hidden in discussions and act like a Loyal Servant to protect yourself (for evil players, Merlin and Percival)
Express that you don’t have any information on whom to put on the team to protect yourself (for Merlin)
Pretend that you don’t have enough information about who is evil right now (for Merlin)
Deception Pretend to be Percival by telling others you are Percival (for evil/Servant)
Pretend to be Merlin by providing hints on who is evil (for Percival and Servant)
Pretend to have information and act like Merlin (for Morgana)
Express you are a loyal player (for evil players)
Express confusion for failure of mission (for evil players)
Persuasion Strongly convince the team leader to choose you on team
Reveal identity by telling others you are Percival to persuade other loyal players(for Percival)
Reveal identity by telling others you are Servant to persuade other loyal players(for Servant)
Teamwork Support team proposal
Support one loyal player (for Merlin)
Support your teammate to be on the quest team (for evil players)
Support your teammate by expressing that he is good (for evil players)
Include loyal players on the quest team as leader (Merlin)
Provide hints to loyal players about evil players (Merlin)
Support loyal players of the previous quest team if the mission failed (for Merlin)
Table 9: Our proposed intention set. Impactful intentions are marked in red.
Structured Context - Intention Summarization
Name: Player1 Role: Merlin Role Details: You are Merlin. These are your goals and characteristics: - Know the identities of evil players. - Subtly guide your team, especially Percival. - Avoid behaviors that expose your role: overly accusing, being too helpful. - Goal(Main Intent): Win without revealing identity. - Player3 and Player4 are evil players Round: 2 Current Leader: Player4 Current Team: Player1, Player2, Player4 Previous Rounds Team Voting: Round 1: Player1 = disagree, Player2 = agree, Player3 = disagree, Player4 = disagree, Player5 = disagree (players didn"t agree on team) Previous Results: Round 1: Team = Player2, Player4, Player3. Result = failure Previous Rounds Summary: Round 1: [[[[summary]]]] Previous Discussions (in the current round): [[[[player name]]]]: [[[[player’s speech]]]] Your thinking: [[[[player’s thinking in the current round]]]] Your speech: [[[[player’s speech in the current round]]]] Summarize your intent from your dialogues in this round. Select multiple intents from the given options that best match your intentions from the current round. Also, provide an explanation of the intents that you showed in the current round. Remember this is private information to you and won’t be shown to other players. Remember that you can select 2-3 intents and don’t use more than 50 words for explanation. Intent options: [[[[intent options]]]]
Table 10: Example of intent summarization prompt with structured context
Structured Context - Intention Guessing
Name: Player2 Role: Servant Role Details: You are a Loyal Servant of Arthur. These are your goals and characteristics: - No special knowledge, rely on discussion and voting. - Contribute to the success of Quests. - Goal(Main Intent): Win by helping complete Quests and protecting Merlin. Round: 2 Speaker Name: Player1 Current Leader: Player4 Current Team: Player1, Player2, Player4 Previous Rounds Team Voting: Round 1: Player1 = disagree, Player2 = agree, Player3 = disagree, Player4 = disagree, Player5 = disagree (players didn"t agree on team) Previous Results: Round 1: Team = Player2, Player4, Player3. Result = failure Previous Rounds Summary: Round 1: [[[[summary]]]] Previous Discussions(in the current round): [[[[player name]]]]: [[[[player’s speech]]]] [[[[First order prompt - Figure 8]]]] Now, here is Player1’s speech: [[[[player’s speech]]]] Select 2-3 intents without modifications that you think Player1 has from the given options based on your guess of their role and speech. Let’s think step by step before making your decisions.
Table 11: Example of intent guessing prompt with structured context
Figure 7: Intention selection prompt.
Figure 8: First order prompt.
Figure 9: Formulation contemplation prompt.
Figure 10: Intention modification prompt.
Figure 11: Refinement contemplation prompt.
Figure 12: Team proposal change prompt.
Figure 13: Quest action prompt.
Figure 14: Introduction for the human annotations and human study.
Figure 15: Explanation of framework.
Figure 16: Explanation of 6 steps in the game.
Figure 17: Intent selection annotation instruction.
Figure 18: Intent selection annotation instruction.
Figure 19: Intent following annotation instruction.
Figure 20: Intent following annotation instruction.
Figure 21: Intent summarization study instruction.
Figure 22: Intent summarization study instruction.
Figure 23: Intent guessing study instruction.